OpenStudy (anonymous):

Solve for F(9) and F(6) F(x)=1/2x+4

OpenStudy (sally):

F(x)=1/2x+4 F(9)=1/2(9)+4 F(9)=4 1/2+4 = 8 1/2 F(x)=1/2x+4 F(6)=1/2(6)+4 F(6)=3+4 = 7

OpenStudy (sally):

if i am not mistaken you substitude x for that F(X) is.

OpenStudy (anonymous):

Thanks!